OpenStudy (anonymous):

can someone help me? Find the product of (4x + 3y)(4x - 3y). A) 16x2 - 24xy + 9y2 B)16x2 - 9y2 C)16x2 + 24xy + 9y2 D) 16x2 + 9y2

OpenStudy (akashdeepdeb):

Alrighty! :D Do you know any algebraic property like this? \[a^2 - b^2 = (a+b)(a-b)\] ? :D

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

May I join ???:)

OpenStudy (anonymous):

yea:)

OpenStudy (anonymous):

answer is B

OpenStudy (akashdeepdeb):

Use that! :D InVeRsElY!! :D \[(4x + 3y)(4x - 3y) = (4x)^2 - (3y)^2 \] Right? Now square them in the parenthesis. :) Understood? :)

OpenStudy (anonymous):

yes, Thank you.

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

@eddy1120:I think learn is better friend :)

OpenStudy (anonymous):

May I say a thing ? @katiebugg1997 @AkashdeepDeb

OpenStudy (anonymous):

(4x+3y)(4x-3y)=(4x)^2+4x*3y-3y*4x-(3y)^2=16x^2-9y^2

OpenStudy (anonymous):

yes you may, @E.ali

OpenStudy (anonymous):

Thank u Modom !:) Look : (a+b)^2=a^2+2ab+b^2 (a-b)^2=a^2-2ab+b^2 a^2-b^2=(a+b)(a-b) (x+a)(x+b)=x^2+(a+b)x+ab and ... They re very important . I think u can proof them for your self :)

OpenStudy (akashdeepdeb):

Yeah Learn them!! :D

OpenStudy (anonymous):

We have more from this :)

OpenStudy (anonymous):

@katiebugg1997:Got them ???

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Got ???

OpenStudy (anonymous):

Yes, thank you. :))

OpenStudy (anonymous):

Your welcome :)